9. The largest mall in Africa by volume is Mall of Africa, South Africa. It measures approximately 631 m long, 707 yards wide, and 110 ft high. Its volume in cubic feet is 4.83 x 108 ft3. Convert the volume to cubic meters. (1 m = 3.28 ft) (4)

A) 1.37 x 107 C) 483 E)  731 x 100

B) 1.7317 x 104 D) 3.141

10. Do the following unit conversion: 1 kilogram = ______ centigram (3)

A) 100 C) 10-5 E) 105

B) 103 D) 103



1
Expert's answer
2021-06-09T08:04:56-0400

9. Since "1 m = 3.28 ft" , then "1 m^3 = (3.28)^3 ft^3". Then "1ft^3 = \dfrac{1}{3.28^3}m^3". Converting the volume to cubic meters, obtain:


"V = 4.83\times 10^8ft^3 = 4.83\times 10^8\cdot \dfrac{1}{3.28^3}m^3 \approx 1.37\times 10^7m"


Thus, the answer is A.


10. Prefix centi- means "10^{-2}". Kilo- means "10^3". Thus:


"1\space kilogram = 1\times 10^3\space gram = 1\times 10^3\times 10^2 \space centigram = 10^5centigram"


Thus, the answer is E.